# Evidence of partonic collectivity in ultra-relativistic heavy-ion collisions with NCQ scaling of radial flow

### v1

@@ 4.3.1 NCQ scaling at $\sqrt{s_{\rm NN}}=200$ $\rm GeV$
Table 1: Scaling parameter, $S_{f}$ values for scaled AMPT Au+Au results (top) and scaled ALICE [10] Pb+Pb results (bottom).
$S_{f}$ Values
System $(\sqrt{s_{\rm NN}})$ Centrality $p_{\rm T}/n_{q}$ $(m_{T}-m_{0})/n_{q}$
Au+Au (200 GeV) 0-20% $0.072$ $0.046$
20-40% $0.085$ $0.066$
10-20% $0.300$ $0.132$
Pb+Pb (5.02 TeV) 30-40% $0.373$ $0.186$
60-70% $0.629$ $0.373$
4.3.2 NCQ scaling at $\sqrt{s_{\rm NN}}=5.02$ $\rm TeV$
@@
We extend the NCQ scaling analysis to LHC energies by compiling the published ALICE $v_{0}(p_{\rm T})$ data for Pb+Pb collisions at $\sqrt{s_{\rm NN}}=5.02$ TeV [10]. Fig. 7 presents the scaled distributions $v_{0}(p_{\rm T})/n_{q}$ as functions of $p_{\rm T}/n_{q}$ (top) and $(m_{\rm T}-m_{0})/n_{q}$ (bottom) for different centrality classes. A clear NCQ scaling is observed for central (10–20%) and semi-central (30–40%) collisions when the transverse kinetic energy per quark, $(m_{\rm T}-m_{0})/n_{q}$, is used as the scaling variable. In contrast, the scaling with $p_{\rm T}/n_{q}$ remains visibly weaker for the same centrality classes. For peripheral (60–70%) collisions, the scaling is found to break down for both variables, indicating a reduced degree of partonic collectivity. The violation of NCQ scaling in peripheral events can be ascribed to the smaller system size and shorter lifetime of the produced medium, which limit the buildup of strong collective dynamics and possibly enhance the relative contribution from non-flow effects and initial-state fluctuations. Conversely, the improved universality of the $(m_{\rm T}-m_{0})/n_{q}$–scaled distributions across hadron species in central collisions supports a scenario in which the observed collectivity is established at the partonic level.
@@
To quantify the degree of NCQ scaling at LHC energies, we apply the same fitting and evaluation procedure defined in Sec. 4.3.1, using the scaling parameter $S_{f}$ [Eqn. (5)]. The extracted $S_{f}$ values, summarized in Tab. 1, show that the scaling is systematically stronger for $(m_{\rm T}-m_{0})/n_{q}$ than for $p_{\rm T}/n_{q}$, and improves toward more central collisions, while being substantially degraded in peripheral events.
5 Summary
In summary, this Letter presents the first systematic study of Number of Constituent Quark (NCQ) scaling of the novel radial-flow observable $v_{0}(p_{\rm T})$. In the absence of published RHIC data, Au+Au collisions at $\sqrt{s_{\rm NN}}=200$ GeV are analyzed using the string-melting configuration of the multi-phase transport (AMPT) model, while Pb+Pb collisions at $\sqrt{s_{\rm NN}}=5.02$ TeV are investigated using published ALICE data [10], with PYTHIA8/Angantyr employed as a non-collective baseline.
Our main observations can be summarized as follows: (i) The $v_{0}(p_{\rm T})$ spectra obtained from AMPT-SM exhibit characteristic signatures of collective behavior, including mass ordering at lower-$p_{\rm T}$ and a baryon–meson separation at intermediate-$p_{\rm T}$, whereas such features are absent in the PYTHIA8/Angantyr calculations. (ii) A direct comparison with ALICE Pb+Pb data reveals that PYTHIA8/Angantyr fails to reproduce the observed $v_{0}(p_{\rm T})$ systematics, providing strong evidence that the experimentally observed correlations cannot be explained without invoking collective dynamics. (iii) Since mass ordering alone is insufficient to establish the microscopic origin of collectivity, we perform a detailed NCQ-scaling analysis. For AMPT-SM, the $v_{0}(p_{\rm T})$ values for pions, kaons, and protons exhibit robust NCQ scaling in central collisions, while clear violations are observed toward peripheral events. (iv) The scaling is found to be significantly improved when expressed in terms of transverse kinetic energy per quark, $(m_{\rm T}-m_{0})/n_{q}$, compared to $p_{\rm T}/n_{q}$, consistent with observations previously reported for elliptic flow $v_{2}$ [36]. (v) A comparative study further indicates that the NCQ scaling of $v_{0}(p_{\rm T})$ is more precise at RHIC energies than at the LHC, following a hierarchy similar to that established for anisotropic flow.
The observation of NCQ scaling in $v_{0}(p_{\rm T})$ provides compelling evidence that the long-range momentum correlations responsible for radial expansion are established at the partonic level of the fireball evolution. This result extends the paradigm of quark-level collectivity from anisotropic flow observables to isotropic radial dynamics, offering a unified and more complete picture of the nearly perfect fluid behavior of the strongly interacting quark-gluon plasma.

### v2

@@ IV.8.1 NCQ scaling at $\sqrt{s_{\rm NN}}=200$ $\rm GeV$
Table 3: Scaling parameter $S_{f}$ for scaled AMPT Au+Au results (top) and scaled ALICE [3] Pb+Pb results (bottom).
System $(\sqrt{s_{\rm NN}})$ Centrality $S_{f}$
$p_{\rm T}/n_{q}$ $(m_{\rm T}-m_{0})/n_{q}$
Au+Au ($200$ GeV) $20\% 0$.0750.075 $0.047$
[…]
IV.8.2 NCQ scaling at $\sqrt{s_{\rm NN}}=5.02$ $\rm TeV$
@@
The NCQ scaling analysis at LHC energy has previously been carried out by the ALICE Collaboration for Pb+Pb collisions at $\sqrt{s_{\rm NN}}=5.02$ TeV [3], though the goodness of the scaling was not rigorously established. Fig. 13 presents the scaled distributions $v_{0}(p_{\rm T})/n_{q}$ as functions of $p_{\rm T}/n_{q}$ (top) and $(m_{\rm T}-m_{0})/n_{q}$ (bottom) for different centrality classes, compiled from the published ALICE data [3]. Only the statistical uncertainties are considered. A clear NCQ scaling is observed for central (10–20%) and semi-central (30–40%) collisions when the transverse kinetic energy per quark, $(m_{\rm T}-m_{0})/n_{q}$, is used as the scaling variable. In contrast, the scaling with $p_{\rm T}/n_{q}$ remains visibly weaker for the same centrality classes. For peripheral (60–70%) collisions, the scaling is found to break down for both variables, indicating a reduced degree of partonic collectivity. The violation of NCQ scaling in peripheral events can be ascribed to the smaller system size and shorter lifetime of the produced medium, which limit the buildup of strong collective dynamics and possibly enhance the relative contribution from non-flow effects and initial-state fluctuations. Conversely, the improved universality of the $(m_{\rm T}-m_{0})/n_{q}$-scaled distributions across hadron species in central collisions supports a scenario in which the observed collectivity is established at the partonic level.
@@
To quantify the degree of NCQ scaling at LHC energies, we apply the same fitting and evaluation procedure defined in Sec. IV.8.1, using the scaling parameter $S_{f}$ [Eq. (7)]. The extracted $S_{f}$ values, summarized in Tab. 3, show that the scaling is systematically stronger for $(m_{\rm T}-m_{0})/n_{q}$ than for $p_{\rm T}/n_{q}$, and improves toward more central collisions, while being substantially degraded in peripheral events.
V Summary
We present a systematic study of radial collectivity in Au+Au collisions at $\sqrt{s_{\rm NN}}=200$ GeV using the differential observable $v_{0}(p_{\rm T})$ within the String Melting configuration of the AMPT model. The $v_{0}(p_{\rm T})$ exhibits a monotonic increase with transverse momentum, evolving from negative values at low-$p_{\rm T}$ to positive values at intermediate $p_{\rm T}$. The sign change occurs around $p_{\rm T}\approx 0.55$$0.63$ GeV/$c$, indicating a characteristic redistribution of spectral yield induced by fluctuations of the radial expansion. The ratio $v_{0}(p_{\rm T})/v_{0}$ demonstrates centrality-independent scaling at low-$p_{\rm T}$, indicating a common dynamical origin of the spectral shape within the partonic stage. The weak dependence of $v_{0}(p_{\rm T})$ on pseudorapidity gap confirms the long-range nature of the correlations, and the observed factorization behavior further supports a collective origin. The $p_{\rm T}$-integrated observable $v_{0}$ decreases toward central collisions and exhibits a sensitivity to the chosen reference momentum interval. However, the normalized quantity $v_{0}/v_{0}^{0\text{--}5\%}$ shows near-universal scaling across different $p_{\rm T}^{\rm ref}$ selections, indicating that $v_{0}$ captures event-by-event fluctuations of the global spectral shape independent of the specific momentum window. Similar collective signatures for inclusive charged hadrons, including long-range pseudorapidity correlations, factorization behavior, and centrality-independent scaling of $v_{0}(p_{\rm T})/v_{0}$, have been reported by the ATLAS Collaboration in Pb+Pb collisions at $\sqrt{s_{\rm NN}}=5.02$ TeV [2], indicating that the global features of radial collectivity observed at LHC energies are also present at RHIC.
For identified particles ($\pi^{\pm},K^{\pm},p+\overline{p}$), $v_{0}(p_{\rm T})$ exhibits mass-ordering at low-$p_{\rm T}$ and meson–baryon separation at intermediate-$p_{\rm T}$. In central collisions, $v_{0}(p_{\rm T})/n_{q}$ follows robust Number of Constituent Quark (NCQ) scaling with $\left(m_{\rm T}-m_{0}\right)/n_{q}$, consistent with partonic collectivity, while this scaling breaks down in peripheral collisions, where the partonic phase of the system becomes increasingly short-lived. The scaling behavior is found to be more precise at $\sqrt{s_{\rm NN}}=200$ GeV than at higher energies, in qualitative agreement with earlier $v_{2}$ measurements [31]. These features are consistent with the observations reported by the ALICE Collaboration in Pb+Pb collisions at $\sqrt{s_{\rm NN}}=5.02$ TeV [3], supporting the interpretation that radial collectivity is predominantly established during the partonic stage and extending the paradigm of quark-level dynamics from anisotropic to isotropic flow.
Appendix A Multiplicity dependence of $v_{0}\sqrt{\rm N_{ch}}$
The relationship between radial flow fluctuations and system size is investigated by examining the product $v_{0}\sqrt{\rm N_{ch}}$ as a function of charged-particle multiplicity (and centrality) in Au+Au collisions at $\sqrt{s_{\rm NN}}=200$ GeV. In Fig. 5, it was seen that $v_{0}$ decreases with increasing multiplicity, approximately following a $1/\sqrt{\rm N_{ch}}$ scaling (verified by the power-law fit, Tab. 2). If this scaling holds, then multiplying $v_{0}$ by $\sqrt{\rm N_{ch}}$ should yield a quantity that remains relatively constant across different centrality classes. Fig. 14 demonstrates this behavior in our AMPT-SM simulations, showing that $v_{0}\sqrt{\rm N_{ch}}$ exhibits only a weak dependence on multiplicity across the centrality range studied. This approximate constancy confirms that the magnitude of global radial flow fluctuations is primarily governed by geometric fluctuations in the initial collision zone, which naturally scale with the inverse square root of the number of participating sources. The mild residual centrality dependence observed, particularly in peripheral collisions, may reflect the increasing role of non-flow correlations or the limitations of the simple independent source picture in describing the full dynamics of the expanding system. These observations are consistent with recent measurements by the ATLAS Collaboration for Pb+Pb collisions [2].
Appendix B Power law fit to the ATLAS $v_{0}$ data
In this appendix, the ATLAS data [2] are replotted and fitted using a power-law. The resulting fit parameters are listed in Tab. 2.
Figure 14: (Color online) $v_{0}\sqrt{\rm N_{ch}}$ versus $\rm N_{ch}$ (left) and centrality (right) for Au+Au collisions at $\sqrt{s_{\rm NN}}=200$ GeV, shown for three $p_{\rm T}^{\rm ref}$ selections.
Figure 15: (Color online) The integrated radial flow measure $v_{0}$ as a function of $\rm N_{ch}$ for three $p_{\rm T}^{\rm ref}$ ranges for Pb+Pb collisions at $\sqrt{s_{\rm NN}}=$ 5.02 TeV. Power-law fits are applied to all $v_{0}$ results up to $\rm N_{ch}^{5\%}$. The figure is compiled using the experimental data of ATLAS [2].